# L-FUSION: Laplacian Fetal Ultrasound Segmentation & Uncertainty Estimation

### v2

@@ 3 Evaluation
Our fetal heart dataset consists of US scans in four-chamber view (4CH), with training data ($n=428$) comprising only normal fetal heart samples, we include $6$ label classes (all chambers, heart and thorax). As OOD samples we test on an example of fetal Congenital Heart Disease (CHD), AtrioVentricular Septal Defect (AVSD) in 4CH. The normal test set includes $92$ normal and $193$ AVSD samples. The public HC18 dataset [9] consists of 1,334 2D US images of fetal standard planes used to measure head circumference (HC), a key metric for assessing gestational age and fetal growth. It includes 999 training images with expert annotations and measurements, and 335 test images without annotations. Images are $800\times 540$ pixels with pixel sizes ranging from 0.052 to 0.326 mm, provided in accompanying metadata files, accessible via Zenodo. As OOD samples we use synthetically generated local deformations on the skull resulting in abnormal skull shape.
Implementation and Training: We extended the $nnj$ library [16], which is used for implementing layers for diagonal backpropagation and fast hessian approximations, by implementing a BatchNorm to reduce exploding gradients and add support for higher-dimensional data (Conv3d, Upsample3d, Batchnorm3d). We resize images to $256\times 256$ for U-Net-based architectures and $224\times 224$ for the UMSF encoder[11]. All evaluated U-Nets have the same architecture besides the Dropout U-Net [12] which has Dropout-layers after each convolutional block. We augment training images with linear transformations and random crop up to $80\%$ of the original size. We align fetal heart scans with an iterative approach, a 2D variant of [22]. All models were trained on an NVIDIA A100 $80$ GB with Early stopping, the number of epochs limited to $250$ for the HC18 dataset and to $400$ (2D) and $600$ (Clip) for the fetal datasets at maximum.
Metrics: Segmentation performance is evaluated using Dice Similarity Coefficient (DSC), Hausdorff Distance (HD), and, Absolute Difference (AD) for HC18 according to the related challenge or Intersection over Union (IoU) for the fetal heart datasets, to quantify overlap and boundary alignment. For out-of-distribution (OOD) performance, we follow [24] and report the Area Under the Receiver Operating Characteristic Curve (AUROC/AUC) across epistemic components, specifically Mutual Information (MI), Expected Pairwise KL (EP), Pixel Variance (PV), and aleatoric uncertainty with Expected Entropy (EE).
Table 1: Results on Head Circumference Dataset HC18; DSC [%] - Dice Similarity Coefficient, HD [mm] - Hausdorff Distance, AD [mm] - Absolute Difference; 1st-ranked, 2nd-ranked.
@@
Quantitive Results: L-FUSION outperforms all baselines on HC18, Tab. 1. It achieves the highest DSC ($95.5~{}\%$) and the lowest HD ($2.7~{}mm$) and AD ($4.7~{}mm$), surpassing Dropout U-Net by $1.3~{}\%$ in DSC and reducing HD by $34.1~{}\%$. Performance gains are consistent across percentiles and the first trimester, where L-FUSION improves DSC by $1.5~{}\%$ over Dropout U-Net while reducing AD by $36.6~{}\%$. Results demonstrate the superiority of L-FUSION in head circumference segmentation. For fetal heart ultrasound (Tab. 2), L-FUSION sets new benchmarks. In the 4CH plane without alignment, it outperforms the best baseline by $12.8\%$ in DSC and reduces HD by $41.6\%$. Under thorax- and heart-aligned conditions, it leads by $9.5\%$ and $0.9\%$ in DSC, respectively, while consistently achieving the lowest HD. L-FUSION also cuts training time by up to $40\%$, with greater savings under stronger alignment, affirming its state-of-the-art status in fetal heart and head circumference segmentation.
@@
Tab. 3 compares the best-performing individual OOD measures for epistemic and aleatoric uncertainty against OOD-Calibration, which scales predicted logit variance ($n=50$) to enhance OOD detection AUC. Applied within the same model (Intra-Model) or across collaborating models (Inter-Model), OOD-Calibration significantly improves AUC. On unaligned 4CH and HC18 datasets, Inter-Model calibration increased AUC by $19\%$ (HC18) and $20\%$ (4CH). For the aligned 4CH dataset, Intra- and Inter-Model calibration performed equally, yielding AUC gains of $3-9\%$.
Table 3: AUC for OOD Detection with OOD-Calibration (using the best-scoring OOD measure for AUC: Expected Entropy, Expected Pairwise Kullback-Leibler, Pixel Variance, Mutal Information); A. - Aligned to Thorax (T) or Heart (H); D. - Dropout, U. - U-Net, L. - Laplace, B - UMSF $+$ D. Segmentation Head, A - UMSF $+$ D.U.; HC18 - Synthetic HC18 datset with Clips;1st-ranked, 2nd-ranked.
OOD - Calibration
Indiv. OOD Measures Intra-Model Inter-Model
Data A. D.U.[12] L.U.[24] L.[11] D.[11] D.U.[12] L.U.[24] L.[11] D.[11] B A
@@
4CH $0.61^{\circ}$ $0.56^{\circ}$ $0.57^{\dagger}$ $0.61^{\circ}$ $0.61^{\circ}$ $0.51^{\circ}$ $0.54^{\ast}$ $0.67^{\ast}$ $\mathbf{0.84}^{\ast}$ $\underline{0.73}^{\ast}$
4CH T $0.61^{\circ}$ $0.54^{\circ}$ $0.56^{\circ}$ $0.56^{\circ}$ $\mathbf{0.89}^{\ast}$ $0.53^{\ast}$ $0.54^{\ast}$ $0.72^{\ast}$ $0.64^{\ast}$ $\underline{0.82}^{\dagger}$
4CH H $0.64^{\diamond}$ $0.59^{\dagger}$ $0.55^{\ast}$ $0.56^{\ast}$ $\mathbf{0.85}^{\circ}$ $0.54^{\ast}$ $0.51^{\circ}$ $0.64^{\circ}$ $0.71^{\ast}$ $\underline{0.82}^{\dagger}$
@@
4CH $\mathbf{0.88^{\dagger}}$ $0.77^{\dagger}$ $0.71^{\ast}$ $\underline{0.81^{\diamond}}$ $\underline{0.81^{\circ}}$ $0.58^{\dagger}$ $0.55^{\circ}$ $\mathbf{0.88^{\circ}}$ $\mathbf{0.88^{\circ}}$ $\underline{0.81^{\circ}}$
@@
Figure 2: Segmentation and uncertainty maps predicted by L-FUSION for Normal and OOD (AVSD) cases in four-chamber view (4-CH) Heart-aligned.
Qualitative Results: Dropout-based uncertainty maps often produce diffuse estimates, see Fig. 1, tending to over-segmentation and missing small details, hence, failing to model fine-grained epistemic uncertainty. The foundation model with a Laplacian segmentation head improves segmentation accuracy and benefits from fine-grained uncertainty estimates but shows high global uncertainty. Furthermore, L-FUSION’s variance maps of logit differences enable counterfactual analysis and improve visual uncertainty maps from both intra- and inter-model variations, see Fig. 2. This provides a more reliable uncertainty quantification, crucial for applications requiring transparent model confidence.
Discussion: Our approach enables efficient segmentation with fine-grained uncertainty quantification benefiting from two uncertainty-modelling approaches, Laplacian approximation and Monte-Carlo Dropout. With OOD-Calibration, we pushed OOD detection AUC up to $+20\%$. However, performance may decline when domain-specific features are essential or data lies on complex, nonlinear manifolds. Uncertainty quantification also depends on embedding quality—loss of local details or high noise can impair segmentation.
@@ 4 Conclusion
Ultrasound segmentation has advanced but still lags behind CT and MRI in uncertainty quantification and consistency[19]. Foundation models [2, 11] improve accuracy but often neglect uncertainty quantification, crucial for clinical use. By combining a foundation model encoder with a Laplacian segmentation head with fast Hessian approximation, a Dropout unit and OOD-Calibration, we reduce complexity while achieving superior performance. Our approach enhances accuracy and generates localised uncertainty maps, improving reliability and clinical applicability over deterministic and Dropout U-Nets.
Acknowledgements: We thank the volunteers and sonographers at St. Thomas’ Hospital London. We gratefully acknowledge financial support from the Wellcome Trust IEH 102431, EPSRC (EP/S022104/1, EP/S013687/1), EPSRC Centre for Medical Engineering [WT 203148/Z/16/Z], the National Institute for Health Research (NIHR) Biomedical Research Centre (BRC) based at Guy’s and St Thomas’ NHS Foundation Trust and King’s College London and supported by the NIHR Clinical Research Facility (CRF) at Guy’s and St Thomas’. This work was also supported by the UK Research and Innovation London Medical Imaging and Artificial Intelligence Centre. We gratefully acknowledge the financial support provided by the German Academic Exchange Service (DAAD) through the Research Stipend (57745342), 2024.

### v4

@@ 3 Evaluation
Datasets. We use two fetal ultrasound datasets: Fetal Heart (4CH): Our dataset includes $428$ normal fetal four-chamber (4CH) scans for training, annotated with six classes (heart chambers, heart, thorax). The test set contains $92$ normal and $193$ abnormal samples with Atrioventricular Septal Defect (AVSD), used for OOD evaluation. HC18 [11]: A public dataset of 1,334 2D fetal head ultrasound images used for head circumference estimation. It includes $999$ annotated training images and $335$ unannotated test images, available via (Zenodo). OOD samples for HC18 are synthetically generated using elastic deformations to simulate local skull shape variations. Additionally, a synthetic variant of the HC18 dataset, denoted HC18, is constructed by simulating frame-by-frame ultrasound sweeps using continuous affine transformations to emulate realistic temporal acquisition dynamics. Each sequence consists of 16 frames.
Implementation and Training. We use USFM [13], pretrained via self-super-vised Masked Image Modelling (MIM) on the 3M-US dataset without labels or annotations. We extended the nnjnnj library [17] with BatchNorm to prevent exploding gradients and support higher-dimensional layers (Conv3d, Upsample3d, BatchNorm3d). Images are resized to $256\times 256$ for U-Nets and $224\times 224$ for the USFM encoder [13]. All U-Nets share the same architecture except Dropout U-Net [14], which adds Dropout after each convolutional block. Training uses linear augmentations and random crops up to $80\%$. Fetal heart scans are aligned via an iterative 2D method [25]. Models trained on an NVIDIA A100 80 GB with early stopping.
Metrics. Segmentation is evaluated using Dice (DSC), Hausdorff Distance (HD), and Absolute Difference (AD; average absolute error) for HC18, and Intersection over Union (IoU) for fetal heart datasets, measuring overlap and boundary accuracy. For OOD detection, following [27], we report AUROC/AUC over epistemic measures—Mutual Information (MI), Expected Pairwise KL (EP), Pixel Variance (PV)—and aleatoric uncertainty via Expected Entropy (EE).
Table 1: Results on Head Circumference Dataset HC18; DSC [%] - Dice Similarity Coefficient, HD [$m$] - Hausdorff Distance, AD [$m$] - Absolute Difference; 1st-ranked, 2nd-ranked.
@@
Quantitive Results. L-FUSION outperforms all baselines on HC18 (Tab. 1), achieving the highest DSC ($95.5\%$) and lowest HD (2.7mm2.7~$m$) and AD (4.7mm4.7~$m$), surpassing Dropout U-Net by $1.3\%$ DSC and reducing HD by $34.1\%$. Gains persist across percentiles and the first trimester, with DSC up $1.5\%$ and AD down $36.6\%$. It also shows the lowest standard deviations and best percentile metrics overall. For fetal heart ultrasound (Tab. 2), L-FUSION sets new benchmarks. In unaligned 4CH, it exceeds the best baseline by $12.8\%$ DSC and cuts HD by $41.6\%$. With thorax- and heart-alignment, it leads by $9.5\%$ and $0.9\%$ DSC, consistently achieving the lowest HD. Training time drops by up to $40\%$, with greater savings under stronger alignment, confirming its state-of-the-art status in fetal heart and head circumference segmentation.
@@
[…]
Table 3: AUC for OOD Detection with Uncertainty-Calibration (identification of best-scoring uncertainty measure $\tilde{U}_{L/D,k}$): Expected Entropy, Expected Pairwise Kullback-Leibler, Pixel Variance, Mutal Information); A. - Aligned to Thorax (T) or Heart (H); D. - Dropout, U. - U-Net, L. - Laplace, (C-1) UMSF ++ D. Segmentation Head, (C-2) UMSF ++ D.U.; 1st-ranked, 2nd-ranked.
Uncertainty - Calibration
Indiv. OOD Measures Intra-Path Inter-Path
Data A. D.U.[14] L.U.[27] L.[13] D.[13] D.U.[14] L.U.[27] L.[13] D.[13] C-1 C-2
@@
4CH $0.61^{\circ}$ $0.56^{\circ}$ $0.57^{\dagger}$ $^{\circ} 0.610.61^{\circ}$ $0.51^{\circ}$ $0.54^{\ast}$ $0.67^{\ast}$ $\mathbf{0.84}^{\ast}$ $\underline{0.73}^{\ast}$
4CH T $0.61^{\circ}$ $0.54^{\circ}$ $^{\circ} 0.560.56^{\circ}$ $\mathbf{0.89}^{\ast}$ $0.53^{\ast}$ $0.54^{\ast}$ $0.72^{\ast}$ $0.64^{\ast}$ $\underline{0.82}^{\dagger}$
4CH H $0.64^{\diamond}$ $9^{\dagger} 0.550$$5^{\ast}$ $0.56^{\ast}$ $\mathbf{0.85}^{\circ}$ $0.54^{\ast}$ $0.51^{\circ}$ $0.64^{\circ}$ $0.71^{\ast}$ $\underline{0.82}^{\dagger}$
@@
4CH $\mathbf{0.88^{\dagger}}$ $0.77^{\dagger}$ $0.71^{\ast}$ $\underline{0.81^{\diamond}}$ $\underline{0.81^{\circ}}$ $8^{\dagger} 0.550$$5^{\circ}$ $\mathbf{0.88^{\circ}}$ $\mathbf{0.88^{\circ}}$ $\underline{0.81^{\circ}}$
@@
Figure 2: Segmentation and Enhanced Uncertainty maps $U_{fused}$ by L-FUSION for Normal and OOD (AVSD) cases in four-chamber view (4-CH) Heart-aligned; the segmented class corresponds to myocardium and valvular structures.Elevated uncertainty levels indicate increased model uncertainty, which may correspond to abnormalities or deviations from the normal training distribution.
Qualitative Results. Dropout-based uncertainty maps are diffuse and over-segmented, missing fine details and limiting epistemic uncertainty capture (Fig. 1). Combining the foundation model with a Laplacian head improves segmentation and uncertainty detail but increases global uncertainty. L-FUSION fuses inter-path logit variance to enable counterfactual analysis and produce refined uncertainty maps (Fig. 2). In both OOD cases, these maps reveal defects in septal formation and valve abnormalities, confirmed by diagnosis.
Discussion. Our method delivers efficient segmentation with fine-grained uncertainty quantification by combining Laplacian approximation and MC Dropout. With Uncertainty-Calibration, OOD detection AUC improves by up to $20\%$. However, performance may drop when domain-specific features dominate or data lies on complex manifolds. Uncertainty estimates also rely on embedding quality; loss of local detail or high noise can degrade segmentation.
@@ 4 Conclusion
Ultrasound segmentation has advanced but still lags behind CT and MRI in uncertainty quantification and consistency [20]. Although foundation models [2, 13] boost accuracy, they often neglect uncertainty, crucial for clinical safety. Our approach, combining a foundation encoder with a Laplacian head using fast Hessian approximation, a Dropout unit, and uncertainty calibration, reduces model complexity while improving performance. It enhances segmentation accuracy and produces localised uncertainty maps, increasing reliability, aiding abnormality detection, and supporting safer clinical decisions. This superior uncertainty quantification surpasses deterministic and Dropout U-Nets, helping sonographers identify low-confidence areas and flag abnormalities for expert review.
{credits}
4.0.1 Acknowledgements
Support was received from the ERC project MIA-NORMAL 101083647, the State of Bavaria (HTA) and DFG 512819079. HPC resources were provided by NHR@FAU of FAU Erlangen-Nürnberg under the NHR project b180dc. NHR@FAU hardware is partially funded by the DFG – 440719683. We thank the volunteers and sonographers at St. Thomas’ Hospital London. We gratefully acknowledge financial support from the Wellcome Trust IEH 102431, EPSRC (EP/S022104/1, EP/S013687/1), EPSRC Centre for Medical Engineering [WT 203148/Z/16/Z], the National Institute for Health Research (NIHR) Biomedical Research Centre (BRC) based at Guy’s and St Thomas’ NHS Foundation Trust and King’s College London and supported by the NIHR Clinical Research Facility (CRF) at Guy’s and St Thomas’. This work was also supported by the UK Research and Innovation AI Centre for Value Based Healthcare.